# Variation of cones of divisors in a family of varieties – Fano type case

### v3

@@ Abstract.
We investigate the relationship between the Fano type property on fibers over a Zariski dense subset and the global Fano type property. We establish the invariance of Néron-Severi spaces, nef cones, effective cones, movable cones, and Mori chamber decompositions for a family of Fano type varieties after a generic finite base change. Additionally, we show the uniform behavior of the minimal model program for this family. These results are applied to the boundedness problem of Fano type varieties.
@@ 1. Introduction
This natural question seems to be rather subtle  [^1] In a recent preprint, [Kim25] confirms Question 1.1 for families of surfaces. , and we can only provide a partial answer.
@@
By the invariance of plurigenera, varieties of general type deform to varieties of general type (see [HMX18]). Theorem 1.2 can be viewed as a weak analogy to this result for Fano type varieties. The proof of this natural theorem, however, requires deep results on complements as developed in [Birkar19]. With this result in mind, to study the deformation of cones of Fano type varieties in the above cases, we can always assume that each fiber is of Fano type after shrinking the base. It turns out that various cones exhibit nice behaviors in this family. As a first step, we show that, after a generic finite base change and a suitable shrinking of the base, the Néron-Severi space of each fiber coincides with the restriction of the Néron-Severi space of the total space.
@@
Assume that $f:X\to T$ is a projective fibration with $(X,\Delta)$ a klt pair for some effective $\mathbb{R}$-divisor $\Delta$ on $X$. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ satisfies
@@
    Up to a generic finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that ${\mathcal{P}}ic_{X_{T_{0}}/T_{0}}\otimes\mathbb{R}$ is a constant sheaf in the Zariski topology.
@@
In the above theorem, “up to a generic finite base change of $T$” means that we take a generic finite morphism $T^{\prime}\to T$ and replace $T$ with $T^{\prime}$. The non-empty open subset $T_{0}$ is then taken inside this new base $T^{\prime}$. Hence, after possibly further shrinking $T_{0}$, we are indeed performing an étale base change of an open subset of the original base.
@@
Let $f:X\to T$ be a projective fibration with $(X,\Delta)$ a klt pair for some effective $\mathbb{R}$-divisor $\Delta$ on $X$. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ satisfies
@@
Then, up to a generic finite base change of $T$, there exist a birational morphism $Y\to X$ and a non-empty open subset $T_{0}\subset T$ such that $Y\to X$ and each fiber $Y_{t}\to X_{t}$ for $t\in T_{0}$ are small $\mathbb{Q}$-factorial modifications.
@@
    Up to a generic finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that the Néron-Severi space $N^{1}(X_{t})$ as well as the cones ${\rm{Eff}}(X_{t})$, ${\rm{Nef}}(X_{t})$ and ${\rm{Mov}}(X_{t})$ are deformation invariant on $T_{0}$. Moreover, the Mori chamber decomposition on ${\rm{Mov}}(X_{t})$ is also deformation invariant on $T_{0}$.
@@
Next, we explore the uniform behavior of the minimal model program (MMP) for a family of varieties. In the case where the fibers are surfaces, the stability of $(-1)$-curves and the simultaneous blowing down of $(-1)$-curves have long been established (see [Kod63, Iit70]). We demonstrate the good behavior of the minimal model program for families of Fano type varieties. As a direct consequence, we obtain the deformation invariance of the Mori chamber decompositions of the movable cones. This circle of ideas has already been applied in [dFH11, HX15]. In the general setting, we establish the following deformation invariance of the MMP.
@@
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, after a generic finite base change of $T$, there exists a non-empty Zariski open subset $T_{0}\subset T$ such that any Zariski open subset $U\subset T_{0}$ satisfies the following properties.
@@
    For any $\mathbb{R}$-divisor $\mathcal{D}\in N^{1}(X_{U}/U)$, a sequence of $\mathcal{D}$-MMP$/U$ induces a sequence of $\mathcal{D}|_{X_{t}}$-MMP of the same type for each $t\in U$.
@@
    Conversely, any $\mathbb{R}$-divisor $D\in N^{1}(X_{t})$ with $t\in U$, a sequence of $D$-MMP on $X_{t}$ is induced by a sequence of $\mathcal{D}$-MMP$/U$ on $X_{U}/U$ of the same type. Moreover, we can choose $\mathcal{D}$ such that $\mathcal{D}|_{X_{t}}\sim_{\mathbb{R}}D$.
@@
The above result is established following the philosophy that the boundedness of varieties follows from birational boundedness together with the finiteness of models. This approach has been successfully applied to many boundedness problems (see [HX15, HMX18, MST20, FHS20], etc.). The proof of Theorem 1.7 is essentially a restatement of the proof of [HX15, Theorem 1.3], with additional clarification on how to obtain a fiber-wise $\mathbb{Q}$-factorial modification.
@@ 2. Preliminaries
A variety means an integral separated scheme of finite type over $\mathbb{C}$. A point of a variety is understood to be a closed point and an open subset of a variety is meant to be a non-empty Zariski open subset, unless explicitly stated otherwise. For a variety $T$, by ”shrinking $T$” we mean replacing $T$ with a Zariski open subset. A subset $S$ of a variety consists of very general points if there exist at most countably many non-empty open subsets $U_{i}$ such that $S=\cap U_{i}$.
@@
the coefficients of $D$ are greater than $-1$. Note that in (2.1), $K_{Y}$ is chosen to be the unique Weil divisor on $Y$ such that $\pi_{*}K_{Y}=K_{X}$. Similarly, if the coefficients of $D$ are greater than or equal to $-1$, then $(X,\Delta)$ is said to have lc singularities. See [KM98, §2.3] for more detailed discussions.
Let $f:X\to T$ be a projective morphism between normal varieties. We use “$/T$” to denote properties which are relative to $T$. Then $X$ is of Fano type$/T$ if there exists a divisor $\Delta$ such that $(X,\Delta)$ has klt singularities and $-(K_{X}+\Delta)$ is ample$/T$. Note that in the definition, we do not assume that $K_{X}$ is a $\mathbb{Q}$-Cartier divisor. By passing to a small $\mathbb{Q}$-factorization, it is straightforward to see that in the definition of Fano type variety, $\Delta$ can be chosen to be a $\mathbb{Q}$-divisor. Moreover, if $K_{X}$ is a $\mathbb{Q}$-Cartier divisor, then $X$ is of Fano type$/T$ if and only if there exists a $\mathbb{Q}$-Cartier divisor $B$ which is big over $T$ such that $(X,B)$ has klt singularities and $K_{X}+B\sim_{\mathbb{Q}}0/T$. Besides, $X\to T$ is called a family of Fano type varieties if all of its closed fibers are Fano type varieties. A variety $X$ is called weak Fano if $-K_{X}$ is nef and big. Hence, a klt weak Fano variety is automatically of Fano type. For some $n\in\mathbb{N}$, a divisor $\Delta\in|-nK_{X}|$ is called an lc (resp. klt) $n$-complement of $K_{X}$ if $(X,\frac{1}{n}\Delta)$ has lc (resp. klt) singularities (see [PS09] for a more general definition). A set of varieties $\mathcal{P}$ is said to admit bounded lc (resp. klt) complements if there exists an integer $n$, depending only on $\mathcal{P}$, such that every element of $\mathcal{P}$ has an lc (resp. klt) $n$-complement.
A Cartier divisor $D$ is movable$/T$ on $X/T$ if the codimension of the support of the sheaf ${\rm coker}(f^{*}f_{*}\mathcal{O}_{X}(D)\to\mathcal{O}_{X}(D))$ is greater than or equal to $2$. We list relevant vector spaces and cones which appear in this paper:
@@
    ${\rm{Pic}}(X/T)$: the $\mathbb{R}$-vector space generated by Cartier divisors on $X$ modulo linear equivalences over $T$, which is the same as ${\rm{Pic}}(X/T)_{\mathbb{Z}}\otimes_{\mathbb{Z}}\mathbb{R}$
@@
If $T={\rm{Spec}}(\mathbb{C})$, then we will omit $T$ in the above notations. For simplicity, when we say that $D\in N^{1}(X/T)$ is a divisor, we mean that $[D]\in N^{1}(X/T)$ with $D$ an $\mathbb{R}$-Cartier divisor. This convention applies to other cones as well. We use ${\rm Int}(C)$ to denote the relative interior of the cone $C$. A cone $C\subset N^{1}(X/T)$ is called a rational polyhedral cone if it is generated by finitely many rational rays (hence, it is a closed cone). In general, ${\rm{Pic}}(X/T)$ may be an infinite dimensional vector space. However, when $X/T$ is of Fano type, it is a finite-dimensional vector space. Besides, many of the above cones are neither open nor closed. Even worse, their closures may not be generated by rational rays. In the literature, there are various conventions regarding effective and movable cones (e.g., taking the closure of the cone or selecting rational elements within the cones). However, when $X/T$ is of Fano type, ${\rm{NE}}(X/T),{\rm{Nef}}(X/T),{\rm{Eff}}(X/T)$ and ${\rm{Mov}}(X/T)$ are all closed rational polyhedral cones. Since this paper focuses on Fano type varieties, we do not need to distinguish these variations on the definitions of cones.
@@ 3. Families with fibers of Fano type varieties over Zariski dense subsets
First, after shrinking $T$, we still have $h({\rm{Exc}}(h))\subset{\rm{Supp}}(\Delta)$. This is because $K_{X}$ is $\mathbb{Q}$-Cartier on $X\backslash{\rm{Supp}}(\Delta)$. As $X_{s},s\in S$ is of Fano type with $S$ a Zariski dense subset, we see that $X\backslash{\rm{Supp}}(\Delta)$ is klt after shrinking $T$. This implies that $h({\rm{Exc}}(h))\subset{\rm{Supp}}(\Delta)$.
@@
and $K_{Z}+\phi_{*}{\rm{Exc}}(h)$ is nef$/X$, we have
$0<(K_{Z}++\phi_{*}{\rm{Exc}}(h))\cdot\ell+\Delta_{Z}\cdot\ell=q^{*}(K_{X}+\Delta)\cdot\ell=0$
@@
First, we claim that there exist a universal $m\in\mathbb{Z}_{>0}$ and a Weil divisor $B^{s}\in|-mK_{X_{s}}|$ such that $(X_{s},\frac{1}{m}B^{s})$ has lc singularities for any $s\in S$. As $X_{s}$ is of Fano type, we can replace it with a small $\mathbb{Q}$-factorial modification $Y_{s}\to X_{s}$. Then $Y_{s}$ is still of Fano type, and it suffices to show the claim for $Y_{s}$. Therefore, we can assume that $X_{s}$ is $\mathbb{Q}$-factorial. Runing a $(-K_{X_{s}})$-MMP, $X_{s}\dasharrow X^{\prime}_{s}$, we have that $X^{\prime}_{s}$ is a Fano type variety with $-K_{X^{\prime}_{s}}$ nef. By [Birkar19, Theorem 1.7], there exist a universal $m\in\mathbb{Z}_{>0}$ and a Weil divisor $D^{s}\in|-mK_{X^{\prime}_{s}}|$ such that $(X^{\prime}_{s},\frac{1}{m}D^{s})$ has lc singularities. Let $p:W\to X_{s},q:W\to X^{\prime}_{s}$ be a resolution of $X_{s}\dasharrow X^{\prime}_{s}$. Then we have
@@
Step 4. In this step, we show that Case (i) can be reduced to Case (iii). As $X$ is $\mathbb{Q}$-factorial, shrinking $T$, there exists $n\in\mathbb{N}$ such that $nK_{X_{s}}$ is Cartier for each $s\in S$. As $-K_{X_{s}}$ is nef and big, by the effective base-point free theorem ([Kol93]), there exists $m\in\mathbb{N}$ such that $|-mK_{X_{s}}|$ is base-point free for each $s\in S$. Since the claim holds when $\dim X_{s}=1$ (i.e., $X_{s}\simeq\mathbb{P}^{1}$), we can assume that $\dim X_{s}\geq 2$. Moreover, without loss of generality, we can also assume that $m\geq 2$. As $-K_{X_{s}}$ is big and $\dim X_{s}\geq 2$, the base-point free linear system $|-mK_{X_{s}}|$ is not composed of a pencil. By Bertini theorem, a general element $B\in|-mK_{X_{s}}|$ is such that the pair $(X,\frac{1}{m}B)$ has klt singularities. This shows that $\{X_{s}\mid s\in S\}$ admits bounded klt complements.
@@
is surjective for each $s\in S$. Therefore, if $\Delta_{s}\in|-nK_{X_{s}}|$ such that $(X_{s},\frac{1}{n}\Delta_{s})$ has klt singularities, then there exists $\Delta\in|-nK_{X}|$ such that $(X,\frac{1}{n}\Delta)$ has klt singularities in a Zariski neighborhood of $X_{s}$ (see [Laz04b, Theorem 9.5.19]). Shrinking $T$, we can assume that $(X,\frac{1}{n}\Delta)$ has klt singularities with $K_{X}+\frac{1}{n}\Delta\sim_{\mathbb{Q}}0$. Hence, there exists $\epsilon\in\mathbb{Q}_{>0}$ such that $(X,(\frac{1}{n}+\epsilon)\Delta)$ still has klt singularities. By [HMX18, Theorem 4.2], after shrinking $T$, we know that the Kodaira dimension $\kappa(X_{t},K_{X_{t}}+(\frac{1}{n}+\epsilon)\Delta_{t})$ is independent of $t\in T$. As $K_{X}+(\frac{1}{n}+\epsilon)\Delta\sim_{\mathbb{Q}}\epsilon n(-K_{X})$, we see that $-K_{X}$ is big over $T$. This implies that $X$ is of Fano type over $T$. ∎

### v4

@@ Abstract.
We investigate the relationship between the Fano type property on fibers over a Zariski dense subset and the global Fano type property. We establish the invariance of Néron-Severi spaces, nef cones, effective cones, movable cones, and Mori chamber decompositions for a family of Fano type varieties after a generically finite base change. Additionally, we show the uniform behavior of the minimal model program for this family. These results are applied to the boundedness problem of Fano type varieties.
@@ 1. Introduction
This natural question seems to be rather subtle, and we can only provide a partial answer.
@@
By the invariance of plurigenera, varieties of general type deform to varieties of general type (see [HMX18]). The above theorem can be viewed as a weak analogy to this result for Fano type varieties. The proof of this natural theorem, however, requires deep results on complements as developed in [Birkar19]. With this result in mind, to study the deformation of cones of Fano type varieties in the above cases, we can always assume that each fiber is of Fano type after shrinking the base. It turns out that various cones exhibit nice behaviors in this family. As a first step, we show that, after a generically finite base change and a suitable shrinking of the base, the Néron-Severi space of each fiber coincides with the restriction of the Néron-Severi space of the total space.
@@
Let $f:X\to T$ be a projective fibration. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ has rational singularities and
@@
    Up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that ${\mathcal{P}}ic_{X_{T_{0}}/T_{0}}\otimes\mathbb{R}$ is a constant sheaf in the Zariski topology.
@@
In the above theorem, “up to a generically finite base change of TT” means that we take a generically finite morphism $T^{\prime}\to T$ and replace $T$ with $T^{\prime}$. The non-empty open subset $T_{0}$ is then taken inside this new base $T^{\prime}$. Hence, after possibly further shrinking $T_{0}$, we are indeed performing an étale base change of an open subset of the original base.
@@
Let $f:X\to T$ be a projective fibration. Suppose that $S\subset T$ is a Zariski dense subset such that for any $s\in S$, the fiber $X_{s}$ has rational singularities and
@@
Assume that there exists a divisor $\Delta$ such that $(X,\Delta)$ has klt singularities. Then, up to a generically finite base change of $T$, there exist a birational morphism $Y\to X$ and a non-empty open subset $T_{0}\subset T$ such that $Y\to X$ and each fiber $Y_{t}\to X_{t}$ for $t\in T_{0}$ are small $\mathbb{Q}$-factorial modifications.
@@
    Up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that the Néron-Severi space $N^{1}(X_{t})$ as well as the cones ${\rm{Eff}}(X_{t})$, ${\rm{Nef}}(X_{t})$ and ${\rm{Mov}}(X_{t})$ are deformation invariant on $T_{0}$. Moreover, the Mori chamber decomposition on ${\rm{Mov}}(X_{t})$ is also deformation invariant on $T_{0}$.
@@
Next, we explore the uniform behavior of the minimal model program (MMP) for a family of varieties. In the case where the fibers are surfaces, the stability of (1)(-1)-curves and the simultaneous blowing down of (1)(-1)-curves have long been established (see [Kod63, Iit70]). We demonstrate the good behavior of the minimal model program for families of Fano type varieties. As a direct consequence, we obtain the deformation invariance of the Mori chamber decompositions of the movable cones. This circle of ideas has already been applied in [dFH11, HX15]. In the general setting, we establish the following deformation invariance of the MMP.
@@
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, after a generically finite base change of $T$, there exists a non-empty Zariski open subset $T_{0}\subset T$ such that any Zariski open subset $U\subset T_{0}$ satisfies the following properties.
@@
    Suppose that $X^{\prime}/U\in{\rm{bc}}(X_{U}/U)$. If $g:X^{\prime}\dashrightarrow Y/U$ is a birational contraction with $Y$ a $\mathbb{Q}$-factorial variety, then for any $t\in U$, the induced map $g_{t}:X^{\prime}_{t}\dashrightarrow Y_{t}$ is again a birational contraction. Moreover, for any $\mathbb{R}$-Cartier divisor $\mathcal{D}$ on $X^{\prime}$, we have
    $(g_{t})_{*}(\mathcal{D}|_{X^{\prime}_{t}})\sim_{\mathbb{R}}(g_{*}\mathcal{D})|_{Y_{t}},\quad t\in U.$
@@
    For any $\mathbb{R}$-divisor $\mathcal{D}\in N^{1}(X_{U}/U)$, a sequence of $\mathcal{D}$-MMP/U/U induces a sequence of $\mathcal{D}|_{X_{t}}$-MMP of the same type for each $t\in U$.
  3. (3)
    Conversely, any $\mathbb{R}$-divisor $D\in N^{1}(X_{t})$ with $t\in U$, a sequence of $D$-MMP on $X_{t}$ is induced by a sequence of $\mathcal{D}$-MMP/U/U on $X_{U}/U$ of the same type. Moreover, we can choose $\mathcal{D}$ such that $\mathcal{D}|_{X_{t}}\sim_{\mathbb{R}}D$.
@@
Theorem 1.7 is proved following the guiding principle that boundedness of varieties can often be deduced from birational boundedness together with the finiteness of models. This strategy has been successfully applied to many boundedness problems (see [HX15, HMX18, MST20, FHS20], etc.). The proof of Theorem 1.7 is essentially a restatement of the argument of [HX15, Theorem 1.3], with additional clarification on how to obtain a fiber-wise $\mathbb{Q}$-factorial modification.
Theorems 1.5, 1.6, and 1.7 are further generalized and applied in the subsequent paper [CLLZ25], where a hybrid of Mori dream spaces and Calabi-Yau type varieties (termed Morrison-Kawamata dream spaces) is investigated.
@@ 2. Preliminaries
A variety means an integral separated scheme of finite type over $\mathbb{C}$. A point of a variety is understood to be a closed point and an open subset of a variety is meant to be a non-empty Zariski open subset, unless explicitly stated otherwise. For a variety $T$, by ”shrinking TT” we mean replacing $T$ with a Zariski open subset. A subset $S$ of a variety consists of very general points if there exist at most countably many non-empty open subsets $U_{i}$ such that $S=\cap U_{i}$.
@@
the coefficients of $D$ are greater than 1-1. Note that in (2.1), $K_{Y}$ is chosen to be the unique Weil divisor on $Y$ such that $\pi_{*}K_{Y}=K_{X}$. Similarly, if the coefficients of $D$ are greater than or equal to 1-1, then $(X,\Delta)$ is said to have lc singularities. See [KM98, §2.3] for more detailed discussions.
Let $f:X\to T$ be a projective morphism between normal varieties. We use “/T/T” to denote properties which are relative to $T$. Then $X$ is of Fano type/T/T if there exists a divisor $\Delta$ such that $(X,\Delta)$ has klt singularities and $(K_{X}+\Delta)$ is ample/T/T. Note that in the definition, we do not assume that $K_{X}$ is a $\mathbb{Q}$-Cartier divisor. By passing to a small $\mathbb{Q}$-factorization, it is straightforward to see that in the definition of Fano type variety, $\Delta$ can be chosen to be a $\mathbb{Q}$-divisor. Moreover, if $K_{X}$ is a $\mathbb{Q}$-Cartier divisor, then $X$ is of Fano type/T/T if and only if there exists a $\mathbb{Q}$-Cartier divisor $B$ which is big over $T$ such that $(X,B)$ has klt singularities and $K_{X}+B\sim_{\mathbb{Q}}0/T$. Besides, $X\to T$ is called a family of Fano type varieties if all of its closed fibers are Fano type varieties. A variety $X$ is called weak Fano if $K_{X}$ is nef and big. Hence, a klt weak Fano variety is automatically of Fano type. For some $n\in\mathbb{N}$, a divisor $\Delta\in|-nK_{X}|$ is called an lc (resp. klt) $n$-complement of $K_{X}$ if $(X,\frac{1}{n}\Delta)$ has lc (resp. klt) singularities (see [PS09] for a more general definition). A set of varieties $\mathcal{P}$ is said to admit bounded lc (resp. klt) complements if there exists an integer $n$, depending only on $\mathcal{P}$, such that every element of $\mathcal{P}$ has an lc (resp. klt) $n$-complement.
A Cartier divisor $D$ is movable/T/T on $X/T$ if the codimension of the support of the sheaf ${\rm coker}(f^{*}f_{*}\mathcal{O}_{X}(D)\to\mathcal{O}_{X}(D))$ is greater than or equal to $2$. We list relevant vector spaces and cones which appear in this paper:
@@
    ${\rm{Pic}}(X/T)$: the $\mathbb{R}$-vector space generated by Cartier divisors on $X$ modulo linear equivalences over $T$, which is the same as ${\rm{Pic}}(X/T)_{\mathbb{Z}}\otimes_{\mathbb{Z}}\mathbb{R}$ .
@@
If $T={\rm{Spec}}(\mathbb{C})$, then we will omit $T$ in the above notations. For simplicity, when we say that $D\in N^{1}(X/T)$ is a divisor, we mean that $D]\in N^{1}(X/T)$ with $D$ an $\mathbb{R}$-Cartier divisor. This convention applies to other cones as well. We use ${\rm Int}(C)$ to denote the relative interior of the cone $C$. A cone $C\subset N^{1}(X/T)$ is called a rational polyhedral cone if it is generated by finitely many rational rays (hence, it is a closed cone). In general, ${\rm{Pic}}(X/T)$ may be an infinite dimensional vector space. However, when $X/T$ is of Fano type, it is a finite-dimensional vector space. Besides, many of the above cones are neither open nor closed. Even worse, their closures may not be generated by rational rays. In the literature, there are various conventions regarding effective and movable cones (e.g., taking the closure of the cone or selecting rational elements within the cones). However, when $X/T$ is of Fano type, ${\rm{NE}}(X/T),{\rm{Nef}}(X/T),{\rm{Eff}}(X/T)$ and ${\rm{Mov}}(X/T)$ are all closed rational polyhedral cones. Since this paper focuses on Fano type varieties, we do not need to distinguish these variations on the definitions of cones.
@@ 3. Families with fibers of Fano type varieties over Zariski dense subsets
[…]
@@
and $K_{Z}+\phi_{*}{\rm{Exc}}(h)$ is nef/X/X, we have
$0<(K_{Z}++\phi_{*}{\rm{Exc}}(h))\cdot\ell+\Delta_{Z}\cdot\ell=q^{*}(K_{X}+\Delta)\cdot\ell=0.$
@@
First, we claim that there exist a universal $m\in\mathbb{Z}_{>0}$ and a Weil divisor $B^{s}\in|-mK_{X_{s}}|$ such that $(X_{s},\frac{1}{m}B^{s})$ has lc singularities for any $s\in S$. As $X_{s}$ is of Fano type, we can replace it with a small $\mathbb{Q}$-factorial modification $Y_{s}\to X_{s}$. Then $Y_{s}$ is still of Fano type, and it suffices to show the claim for $Y_{s}$. Therefore, we can assume that $X_{s}$ is $\mathbb{Q}$-factorial. Runing a $(-K_{X_{s}})$-MMP, $X_{s}\dasharrow X^{\prime}_{s}$, we have that $X^{\prime}_{s}$ is a Fano type variety with $K_{X^{\prime}_{s}}$ nef. By [Birkar19, Theorem 1.7], there exist a universal $m\in\mathbb{Z}_{>0}$ and a Weil divisor $D^{s}\in|-mK_{X^{\prime}_{s}}|$ such that $(X^{\prime}_{s},\frac{1}{m}D^{s})$ has lc singularities. Let $p:W\to X_{s},q:W\to X^{\prime}_{s}$ be a resolution of $X_{s}\dasharrow X^{\prime}_{s}$. Then we have
@@
Step 4. In this step, we show that Case (i) can be reduced to Case (iii). As $X$ is $\mathbb{Q}$-factorial, shrinking $T$, there exists $n\in\mathbb{N}$ such that $nK_{X_{s}}$ is Cartier for each $s\in S$. As $K_{X_{s}}$ is nef and big, by the effective base-point free theorem ([Kol93]), there exists $m\in\mathbb{N}$ such that $|-mK_{X_{s}}|$ is base-point free for each $s\in S$. Since the claim holds when $\dim X_{s}=1$ (i.e., $X_{s}\simeq\mathbb{P}^{1}$), we can assume that $\dim X_{s}\geq 2$. Moreover, without loss of generality, we can also assume that $m\geq 2$. As $K_{X_{s}}$ is big and $\dim X_{s}\geq 2$, the base-point free linear system $|-mK_{X_{s}}|$ is not composed of a pencil. By Bertini theorem, a general element $B\in|-mK_{X_{s}}|$ is such that the pair $(X,\frac{1}{m}B)$ has klt singularities. This shows that $\{X_{s}\mid s\in S\}$ admits bounded klt complements.
@@
is surjective for each $s\in S$. Therefore, if $\Delta_{s}\in|-nK_{X_{s}}|$ such that $(X_{s},\frac{1}{n}\Delta_{s})$ has klt singularities, then there exists $\Delta\in|-nK_{X}|$ such that $(X,\frac{1}{n}\Delta)$ has klt singularities in a Zariski neighborhood of $X_{s}$ (see [Laz04b, Theorem 9.5.19]). Shrinking $T$, we can assume that $(X,\frac{1}{n}\Delta)$ has klt singularities with $K_{X}+\frac{1}{n}\Delta\sim_{\mathbb{Q}}0$. Hence, there exists $\epsilon\in\mathbb{Q}_{>0}$ such that $(X,(\frac{1}{n}+\epsilon)\Delta)$ still has klt singularities. By [HMX18, Theorem 4.2], after shrinking $T$, we know that the Kodaira dimension $\kappa(X_{t},K_{X_{t}}+(\frac{1}{n}+\epsilon)\Delta_{t})$ is independent of $t\in T$. As $K_{X}+(\frac{1}{n}+\epsilon)\Delta\sim_{\mathbb{Q}}\epsilon n(-K_{X})$, we see that $K_{X}$ is big over $T$. This implies that $X$ is of Fano type over $T$. ∎